a sterilization tank contains 120 gallons of a solution of 6% bleach mixed with pure distilled water. for complete sterilization, the solution only needs to be at 4%. how much of the solution needs to be drained and replaced with water to reduce the concentration of bleach.
note:: u want to still have 120 gallons of solution when you are finished diluting.

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

a sterilization tank contains 120 gallons of a solution of 6% bleach mixed with pure distilled water. for complete sterilization, the solution only needs to be at 4%. how much of the solution needs to be drained and replaced with water to reduce the concentration of bleach.
note:: u want to still have 120 gallons of solution when you are finished diluting.
-----------------------
Equation:
bleach - bleach + bleach = bleach
0.06*120 - 0.06x + 0x = 0.04*120
---
Multiply thru by 100 to get:
6*120 - 6x + 0 = 4*120
6x = 2*120
x = 40 gallons (amt. of 6% solution that must be replaced by water)
==========================
Cheers,
Stan H.

Answer by mananth(16946)

sterilization tank contains 120 gallons of a solution of 6% bleach
mixed with pure distilled water.
needs to be at 4%.
..
let water required be x 0%
120-x gallons 6%
Mix 120 gallons 4%
...
0+0.06(120-x)=120*0.04
7.20-0.06x=4.80
7.20-4.80 = 0.06x
2.4 = 0.06x
x=40 liters